# LargeAD: Large-Scale Cross-Sensor Data Pretraining for Autonomous Driving

### v2

@@ 5.1 Datasets
TABLE VI: Comparisons of state-of-the-art pretraining methods [30, 119, 29] for the panoptic LiDAR segmentation task on the Panoptic-nuScenes [123] dataset. All scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
TABLE VII: Comparisons of zero-shot LiDAR segmentation performance on the validation set of nuScenes [123]. All mIoU and mAcc scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
TABLE VIII: Ablation study on the knowledge transfer effects from the classic SLIC [32] algorithm and different vision foundation models [5, 6, 28, 7]. Both SLidR [30] and our framework are pretrained on nuScenes [123] and fine-tuned on nuScenes [113], SemanticKITTI [114], Waymo Open [115], SemanticPOSS [131], SemanticSTF [140], and Synth4D [70], respectively, and evaluated on the official validation split of each dataset. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
TABLE IX: Ablation study of each component in our framework. All variants are pretrained on nuScenes [123] and fine-tuned on nuScenes [123], SemanticKITTI [114], and Waymo Open [115], respectively. The knowledge is distilled from the ResNet-50. C2L: Camera-to-LiDAR distillation. VFM: Vision foundation models. STC: Superpoint temporal consistency. P2S: Point-to-segment regularization. CDP: Cross-dataset pretraining. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
TABLE X: Ablation study on the utilization of multiple data sources during pretraining. The knowledge is distilled from the ViT-B. N, K, and W denote pretraining on nuScenes [123], SemanticKITTI [114], and Waymo Open [115], respectively. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
TABLE XI: Ablation study on the possible sensor misalignment between the LiDAR and camera sensors. The perturbation is randomly generated and inserted during the pretraining stage. All mIoU scores are given in percentage (%).
Method 1% Misalignment 5% Misalignment 10% Misalignment
CL 1% 5% 10% 25% CL 1% 5% 10% 25% CL 1% 5% 10% 25%
5.2 Implementation Details
Evaluation Metrics. For linear probing and downstream tasks on LiDAR semantic segmentation, we report the Intersection-over-Union (IoU) for each semantic class and the mean IoU (mIoU) across all classes. For 3D object detection, we report the mean Average Precision (mAP) and nuScenes Detection Score (NDS). For panoptic LiDAR segmentation, we follow the Panoptic-nuScenes protocol [123] and report the Panoptic Quality (PQ), Recognition Quality (RQ), and Segmentation Quality (SQ). For robustness evaluation, we adopt the Robo3D benchmark [135] and report the mean Corruption Error (mCE) and mean Resilience Rate (mRR) scores. The baseline model used for these evaluations is the MinkUNet implementation by Tang et al.[54].
5.3 Comparative Study
Comparisons to State-of-the-Arts. We compare the proposed LargeAD against random initialization and a total of eleven state-of-the-art pretraining techniques using both linear probing (LP) and few-shot fine-tuning protocols on nuScenes [123], as presented in Tab. I. The results demonstrate the significant impact of pretraining on downstream task performance, particularly in low-data regimes such as $1\%$, $5\%$, and $10\%$ fine-tuning budgets. When distilling the knowledge from ResNet, ViT-S, ViT-B, and ViT-L, our framework achieves mIoU scores of $46.13\%$, $46.58\%$, $47.84\%$, and $48.71\%$ under the LP setting, outperforming the previous best models [31, 29, 127, 129] by large margins. Moreover, our framework consistently delivers the highest performance across almost all fine-tuning tasks on nuScenes, highlighting the effectiveness of VFM-assisted contrastive learning, the incorporation of spatial-temporal consistency regularization, and the combination of multi-source data pretraining.
Downstream Generalization. To thoroughly evaluate the generalization capability of our method, we perform experiments across a total of nine autonomous driving datasets, with the results summarized in Tab. I (for SemanticKITTI and Waymo Open) and Tab. III (for the other seven datasets). Each dataset presents distinct challenges, including variations in sensor types, acquisition environments, scale, and data fidelity, making this a rigorous assessment of model generalization. Our framework achieves $51.68\%$ and $52.68\%$ mIoU scores on SemanticKITTI and Waymo Open when distilling from ViT-L, setting new states for these benchmarks. We also surpass SLidR [30] and Seal [29] on the other seven datasets, as shown in Tab. III. The results consistently show that the proposed method outperforms existing state-of-the-art approaches on all evaluated datasets. These outcomes underline the robustness and adaptability of our proposed approach to a wide range of real-world automotive perception tasks.
Robustness Probing. Assessing the robustness of learned representations on out-of-distribution data is crucial, particularly for real-world applications where environments are unpredictable. We utilize the nuScenes-C dataset from the Robo3D benchmark [135] to evaluate robustness under various corruptions. As shown in Tab. IV, self-supervised learning methods like PPKT [119] and SLidR [30] generally demonstrate better resilience compared to traditional baselines (with the random initialization) such as MinkUNet [49]. Our approach, LargeAD, achieves superior robustness across nearly all corruption types, outperforming other recent segmentation backbones that rely on different LiDAR representations, including range view [44], bird’s eye view (BEV) [47], raw point-based methods [39], and multi-view fusion [54]. These results underscore the adaptability and resilience of our pretraining framework under diverse real-world autonomous driving conditions.
Enhancements for 3D Object Detection. In addition to LiDAR semantic segmentation, we extend our framework to 3D object detection tasks on the nuScenes dataset [113] and compare it with state-of-the-art pretraining methods. The results, shown in Tab. V, indicate that LargeAD consistently outperforms competing approaches across various data proportions ($5\%$, $10\%$, and $20\%$) for both the CenterPoint [136] and SECOND [137] backbones. In particular, our method achieves the highest mAP and NDS at all fine-tuning levels, surpassing recent techniques such as CSC [127] and TriCC [126]. Notably, our framework maintains superior performance with limited fine-tuning data, demonstrating its robustness and effectiveness for 3D object detection. These results further validate the generalizability of our framework across multiple challenging tasks in autonomous driving, from semantic segmentation to object detection.
Enhancements for Panoptic LiDAR Segmentation. To further showcase the general-purpose capability of LargeAD, we evaluate its performance on the challenging panoptic LiDAR segmentation task using the Panoptic-nuScenes benchmark [123]. This task jointly requires semantic-level understanding and instance-level separation, making it a strong indicator of representation richness. As shown in Tab. VI, our model achieves the best results across all metrics (PQ, RQ, and SQ) under both 1% and 5% fine-tuning regimes. Compared to previous methods like SLidR [30] and Seal [29], our framework demonstrates consistent advantages, particularly in instance-level recognition (RQ). These results confirm that the representations learned can effectively support both category discrimination and fine-grained object delineation, further validating our general applicability.
Qualitative Assessment. To further assess the performance of our framework, we visualize the segmentation predictions on nuScenes [123] in Fig. 6. The pretraining methods clearly enhance segmentation quality compared to models trained from random initialization. Among the compared approaches, LargeAD demonstrates the most consistent and accurate results, particularly in complex driving environments. This improvement can be attributed to the robust spatial and temporal consistency learning embedded in our pretraining strategy. Additionally, we perform qualitative comparisons of our framework with the random initialization on the 3D object detection. Qualitative results from Fig. 7 further highlight our superiority. Our framework generates significantly more accurate predictions, closely matching ground truth boxes and outperforming models trained from random initialization, which exhibit greater inaccuracies. The strong performance across both tasks underscores the generalizability and robustness of our pretraining approach.
Extension to Zero-Shot LiDAR Segmentation. To further validate the generalization ability of our pretraining framework, we conduct zero-shot LiDAR semantic segmentation experiments following the protocol of OpenScene [139]. This task bypasses any label supervision by leveraging CLIP-based image-to-point alignment, where semantic predictions are made via cosine similarity with textual category prompts. As shown in Tab. VII, LargeAD achieves superior performance over existing methods, obtaining $43.8\%$ mIoU and $62.7\%$ mAcc on the nuScenes validation set. These results not only highlight the rich cross-modal semantics captured by our pretraining pipeline, but also reinforce its strong transferability in label-scarce or fully unsupervised deployment scenarios.
@@ 5.3 Comparative Study
Figure 9: Temporal failure cases segmented by the RANSAC [121] algorithm. In the first row, dynamic objects are segmented into inconsistent instances across different timestamps, indicating a lack of temporal coherence. In the second row, the moving objects are either incorrectly merged with surrounding background points or fragmented into multiple inconsistent instances due to motion and scene clutter.
5.4 Ablation Study
Comparing Different Foundation Models. This study serves as the first attempt to adapt VFMs for large-scale point cloud representation learning. We conduct a comprehensive ablation study on four popular VFMs, i.e., SAM [5], X-Decoder [6], OpenSeeD [28], and SEEM [7], and show the results in Tab. VIII. Our experiments reveal that different VFMs have varying levels of impact on contrastive learning objectives. All VFMs consistently outperform traditional SLIC [32] when integrated into both frameworks. SEEM [7] emerges as the top performer, providing the most consistent improvements. We attribute this to its ability to produce compact, semantically aligned superpixels that better match LiDAR segments. Interestingly, SAM [5] generates more fine-grained superpixels, which enhances performance when fine-tuning with larger labeled datasets. However, its finer segmentation may lead to redundant or overly fragmented contrastive pairs, explaining the lower performance under limited supervision. We hypothesize that SAM [5] offers more diverse negative samples, which may benefit the superpixel-driven contrastive learning. Across all configurations, our framework surpasses SLidR [30] by a significant margin, affirming the effectiveness of our proposed large-scale cross-sensor data pretraining framework.
Cosine Similarity. We visualize feature similarities across various VFMs in Fig. 5, providing insights into the distinction in representations even before fine-tuning. Semantically enriched models like X-Decoder [6], OpenSeeD [28], and SEEM [7] exhibit clear feature distinctions between objects and backgrounds. In contrast, unsupervised or overly fine-grained methods such as SLIC [32] and SAM [5] show weaker semantic awareness. These qualitative observations are mirrored in the performance results from both linear probing and fine-tuning tasks (see Tab. VIII), where SEEM demonstrates stronger consistency regularization during cross-sensor representation learning, leading to improved downstream task performance.
Component Analysis. The ablation results of our core components are in Tab. IX. Integrating VFMs alone (row c) delivers a $4.20\%$ mIoU improvement in linear probing, while adding temporal consistency learning (row b) yields an additional $1.65\%$ mIoU gain. Combining these two components (row d) results in a total of $5.21\%$ mIoU boost. The point-to-segment regularization (row e) also contributes a significant improvement of $4.55\%$ mIoU on its own. Moreover, when both STC and P2S are used jointly (row f), we observe additional gains beyond their individual contributions, suggesting that the two modules offer complementary benefits. When all components are integrated (row g), the final model achieves a total $6.33\%$ mIoU gain over SLidR [30], outperforming all state-of-the-art methods on both in-distribution and out-of-distribution benchmarks.
Scaling with Data Sources. We conduct an ablation study to examine the impact of using different datasets during pretraining, as summarized in Tab. X. The results demonstrate that pretraining on a single dataset, i.e., nuScenes (N) [123], SemanticKITTI (K) [114], or Waymo Open (W) [115], provides notable improvements over random initialization, especially in the linear probing (LP) and $1\%$ fine-tuning evaluations. However, as more datasets are combined during pretraining, the performance continues to improve consistently across both in-domain (the pretrained dataset) and out-of-domain datasets. For example, pretraining on all three datasets (N + K + W) results in the best performance, achieving a significant boost across all scenarios. Interestingly, the benefits of multi-dataset pretraining are most evident in the out-of-domain results, where combining two or three datasets leads to substantial gains over single-dataset pretraining. For instance, the inclusion of both nuScenes and Waymo Open (N + W) leads to a $47.42\%$ mIoU in LP on nuScenes, outperforming the single-dataset pretraining setups. Similarly, using all three datasets outperforms two-dataset combinations in both LP and fine-tuning, particularly in out-of-domain scenarios like Waymo Open, where it achieves a remarkable $51.52\%$ mIoU in $1\%$ fine-tuning. These results highlight the importance of multi-source pretraining, which not only improves generalization within in-domain datasets but also significantly enhances out-of-domain performance, showcasing the robustness and scalability of our proposed framework.
@@ 5.4 Ablation Study
Robustness to Sensor Misalignment. To evaluate the robustness of our framework to calibration errors between LiDAR and cameras, we simulate misalignment by perturbing the extrinsic transformation matrix $\Gamma_{K}$ with random translation and rotation noise at three levels: 1%, 5%, and 10%. As in Tab. XI, our approach exhibits strong resilience under misalignment, with minimal performance degradation even at 10% perturbation. This robustness can be attributed to the region-level supervision provided by VFM–generated superpixels, which are less sensitive to precise point-to-pixel alignment, and to the superpoint temporal consistency module, which operates entirely in 3D space. In contrast, point-wise methods such as PPKT [119] suffer larger performance drops due to their reliance on accurate 2D–3D correspondences. These results suggest that our framework remains effective even under mild to moderate calibration noise, enhancing its applicability in real-world deployments.
@@
Visual Examples of Failure Cases. In addition to the above discussions, we have included the corresponding visual examples in Fig. 8 and Fig. 9 for further reference. Specifically, we show superpixel failures from different VFMs under occlusion, nighttime, and adverse weather scenes, as well as inconsistent temporal segmentation across frames. These examples offer additional insight into the failure modes of the proposed modules and complement previous discussions.
@@ 6 Conclusion
Acknowledgments. This work is supported by Shanghai Artificial Intelligence Laboratory. This work is affiliated to the PJ-PRJ24DMX001 program.
This work is also supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221-0012, MOE-T2EP20223-0002). This research is also supported by cash and in-kind funding from NTU S-Lab and industry partner(s). Lingdong Kong is supported by the Apple Scholars in AI/ML Ph.D. Fellowship program.

### v3

@@ 5.3 Comparative Study
Robustness Probing. Assessing the robustness of learned representations on out-of-distribution data is crucial, particularly for real-world applications where environments are unpredictable. We utilize the nuScenes-C dataset from the Robo3D benchmark [135] to evaluate robustness under various corruptions. As shown in Table 4, self-supervised learning methods like PPKT [119] and SLidR [29] generally demonstrate better resilience compared to traditional baselines (with the random initialization) such as MinkUNet [48]. Our approach, LargeAD, achieves superior robustness across nearly all corruption types, outperforming other recent segmentation backbones that rely on different LiDAR representations, including range view [43], bird’s eye view (BEV) [46], raw point-based methods [38], and multi-view fusion [53]. These results underscore the adaptability and resilience of our pretraining framework under diverse real-world autonomous driving conditions.
Enhancements for 3D Object Detection. In addition to LiDAR semantic segmentation, we extend our framework to 3D object detection tasks on the nuScenes dataset [113] and compare it with state-of-the-art pretraining methods. The results, shown in Table 5, indicate that LargeAD consistently outperforms competing approaches across various data proportions ($5\%$, $10\%$, and $20\%$) for both the CenterPoint [137] and SECOND [138] backbones. In particular, our method achieves the highest mAP and NDS at all fine-tuning levels, surpassing recent techniques such as CSC [127] and TriCC [126]. Notably, our framework maintains superior performance with limited fine-tuning data, demonstrating its robustness and effectiveness for 3D object detection. These results further validate the generalizability of our framework across multiple challenging tasks in autonomous driving, from semantic segmentation to object detection.
Table 6: Comparisons of state-of-the-art pretraining methods [29, 119, 31] for the panoptic LiDAR segmentation task on the Panoptic-nuScenes [123] dataset. All scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
Enhancements for Panoptic LiDAR Segmentation. To further showcase the general-purpose capability of LargeAD, we evaluate its performance on the panoptic LiDAR segmentation task using the Panoptic-nuScenes benchmark [123]. This task jointly requires semantic-level understanding and instance-level separation, making it a strong indicator of representation richness. As shown in Table 6, our model achieves the best results across all metrics (PQ, RQ, and SQ) under both 1% and 5% fine-tuning regimes. Compared to previous methods like SLidR [29] and Seal [31], our framework demonstrates consistent advantages, particularly in instance-level recognition (RQ). These results confirm that the representations learned can effectively support both category discrimination and fine-grained object delineation, further validating our general applicability.
Qualitative Assessment. To further assess the performance of our framework, we visualize the segmentation predictions on nuScenes [123] in Figure 6. The pretraining methods clearly enhance segmentation quality compared to models trained from random initialization. Among the compared approaches, LargeAD demonstrates the most consistent and accurate results, particularly in complex driving environments. This improvement can be attributed to the robust spatial and temporal consistency learning embedded in our pretraining strategy. Additionally, we perform qualitative comparisons of our framework with the random initialization on the 3D object detection. Qualitative results from Figure 7 further highlight our superiority. Our framework generates significantly more accurate predictions, closely matching ground truth boxes and outperforming models trained from random initialization, which exhibit greater inaccuracies. The strong performance across both tasks underscores the generalizability and robustness of our pretraining approach.
Table 7: Comparisons of zero-shot LiDAR segmentation performance on nuScenes [123]. All mIoU and mAcc scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
Extension to Zero-Shot LiDAR Segmentation. To further validate the generalization ability of our pretraining framework, we conduct zero-shot LiDAR semantic segmentation experiments following the protocol of OpenScene [140]. This task bypasses any label supervision by leveraging CLIP-based image-to-point alignment, where semantic predictions are made via cosine similarity with textual category prompts. As shown in Table 7, LargeAD achieves superior performance over existing methods, obtaining $43.8\%$ mIoU and $62.7\%$ mAcc on the nuScenes validation set. These results not only highlight the rich cross-modal semantics captured by our pretraining pipeline, but also reinforce its strong transferability in label-scarce or fully unsupervised deployment scenarios.
5.4 Ablation Study
Comparing Different Foundation Models. This study serves as the first attempt of adapting VFMs for large-scale point cloud representation learning. We conduct a comprehensive ablation study on four popular VFMs, i.e., SAM [6], X-Decoder [7], OpenSeeD [28], and SEEM [8], and show the results in Table 8. Our experiments reveal that different VFMs have varying levels of impact on contrastive learning objectives. All VFMs consistently outperform traditional SLIC [32] when integrated into both frameworks. SEEM [8] emerges as the top performer, providing the most consistent improvements. We attribute this to its ability to produce compact, semantically aligned superpixels that better match LiDAR segments. Interestingly, SAM [6] generates more fine-grained superpixels, which enhances performance when fine-tuning with larger labeled datasets. However, its finer segmentation may lead to redundant or overly fragmented contrastive pairs, explaining the lower performance under limited supervision. We hypothesize that SAM [6] offers more diverse negative samples, which may benefit the superpixel-driven contrastive learning. Across all configurations, our framework surpasses SLidR [29] by a significant margin, affirming the effectiveness of our proposed large-scale cross-sensor data pretraining framework.
Table 8: Ablation study on knowledge transfer effects from the classic SLIC [32] algorithm and different vision foundation models [6, 7, 28, 8]. Both SLidR [29] and our framework are pretrained on nuScenes [123] and fine-tuned on nuScenes [113], SemanticKITTI [114], Waymo Open [115], SemanticPOSS [131], SemanticSTF [141], and Synth4D [69], respectively, and evaluated on the official validation split of each dataset. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@ 5.4 Ablation Study
Table 9: Ablation study of components in our framework. All variants are pretrained on nuScenes [123] and fine-tuned on nuScenes [123], SemanticKITTI [114], and Waymo [115], respectively. The knowledge is distilled from the ResNet-50. C2L: Camera-to-LiDAR distillation. VFM: Vision foundation models. STC: Superpoint temporal consistency. P2S: Point-to-segment regularization. CDP: Cross-dataset pretraining. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
Table 10: Ablation study on the possible sensor misalignment between the LiDAR and camera sensors. The perturbation is randomly generated and inserted during the pretraining stage. All mIoU scores are given in percentage (%).
Method 1% Misalignment 5% Misalignment 10% Misalignment
CL 1% 5% 10% 25% CL 1% 5% 10% 25% CL 1% 5% 10% 25%
Cosine Similarity. We visualize feature similarities across various VFMs in Figure 5, providing insights into the distinction in representations even before fine-tuning. Semantically enriched models like X-Decoder [7], OpenSeeD [28], and SEEM [8] exhibit clear feature distinctions between objects and backgrounds. In contrast, unsupervised or overly fine-grained methods such as SLIC [32] and SAM [6] show weaker semantic awareness. These qualitative observations are mirrored in the performance results from both linear probing and fine-tuning tasks (see Table 8), where SEEM demonstrates stronger consistency regularization during cross-sensor representation learning, leading to improved downstream task performance.
Component Analysis. The ablation results of our core components are in Table 9. Integrating VFMs alone (row c) delivers a $4.20\%$ mIoU improvement in linear probing, while adding temporal consistency learning (row b) yields an additional $1.65\%$ mIoU gain. Combining these two components (row d) results in a total of $5.21\%$ mIoU boost. The point-to-segment regularization (row e) also contributes a significant improvement of $4.55\%$ mIoU on its own. Moreover, when both STC and P2S are used jointly (row f), we observe additional gains beyond their individual contributions, suggesting that the two modules offer complementary benefits. When all components are integrated (row g), the final model achieves a total $6.33\%$ mIoU gain over SLidR [29], outperforming all state-of-the-art methods on both in-distribution and out-of-distribution benchmarks.
Table 11: Ablation study on the utilization of multiple data sources during pretraining. The knowledge is distilled from the ViT-B. N, K, and W denote pretraining on nuScenes [123], SemanticKITTI [114], and Waymo Open [115], respectively. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
Scaling with Data Sources. We conduct an ablation study to examine the impact of using different datasets during pretraining, as summarized in Table 11. The results demonstrate that pretraining on a single dataset, i.e., nuScenes (N) [123], SemanticKITTI (K) [114], or Waymo Open (W) [115], provides notable improvements over random initialization, especially in the linear probing (LP) and $1\%$ fine-tuning evaluations. However, as more datasets are combined during pretraining, the performance continues to improve consistently across both in-domain (the pretrained dataset) and out-of-domain datasets. For example, pretraining on all three datasets (N + K + W) results in the best performance, achieving a significant boost across all scenarios. Interestingly, the benefits of multi-dataset pretraining are most evident in the out-of-domain results, where combining two or three datasets leads to substantial gains over single-dataset pretraining. For instance, the inclusion of both nuScenes and Waymo Open (N + W) leads to a $47.42\%$ mIoU in LP on nuScenes, outperforming the single-dataset pretraining setups. Similarly, using all three datasets outperforms two-dataset combinations in both LP and fine-tuning, particularly in out-of-domain scenarios like Waymo Open, where it achieves a remarkable $51.52\%$ mIoU in $1\%$ fine-tuning. These results highlight the importance of multi-source pretraining, which not only improves generalization within in-domain datasets but also significantly enhances out-of-domain performance, showcasing the robustness and scalability of our proposed framework.
@@
Robustness to Sensor Misalignment. To evaluate the robustness of our framework to calibration errors between LiDAR and cameras, we simulate misalignment by perturbing the extrinsic transformation matrix $\Gamma_{K}$ with random translation and rotation noise at three levels: 1%, 5%, and 10%. As in Table 10, our approach exhibits strong resilience under misalignment, with minimal performance degradation even at 10% perturbation. This robustness can be attributed to the region-level supervision provided by VFM–generated superpixels, which are less sensitive to precise point-to-pixel alignment, and to the superpoint temporal consistency module, which operates entirely in 3D space. In contrast, point-wise methods such as PPKT [119] suffer larger performance drops due to their reliance on accurate 2D–3D correspondences. These results suggest that our framework remains effective even under mild to moderate calibration noise, enhancing its applicability in real-world deployments.
Figure 9: Temporal failure cases segmented by RANSAC [121]. In the first row, dynamic objects are segmented into inconsistent instances across different timestamps, indicating a lack of temporal coherence. In the second row, the moving objects are either incorrectly merged with surrounding background points or fragmented into multiple inconsistent instances due to motion and scene clutter.
@@
Visual Examples of Failure Cases. In addition to the above discussions, we have included the corresponding visual examples in Figure 8 and Figure 9 for further reference. Specifically, we show superpixel failures from different VFMs under occlusion, nighttime, and adverse weather scenes, as well as inconsistent temporal segmentation across frames. These examples offer additional insight into the failure modes of the proposed modules and complement previous discussions.